# Temporally Coherent Imitation Learning via Latent Action Flow Matching for Robotic Manipulation

### v1

@@ Abstract
Learning long-horizon robotic manipulation requires jointly achieving expressive behavior modeling, real-time inference, and stable execution, which remains challenging for existing generative policies. Diffusion-based approaches offer strong modeling capacity but incur high inference latency, while flow matching enables fast one-step generation yet often suffers from unstable execution when operating directly in the raw action space. We propose LG-Flow Policy, a trajectory-level imitation learning framework that performs flow matching in a continuous latent action space. By encoding action sequences into temporally regularized latent trajectories and learning an explicit latent-space flow, LG-Flow Policy decouples global motion structure from low-level control noise, enabling smooth and reliable long-horizon execution. The framework further integrates geometry-aware point cloud conditioning and execution-time multimodal modulation, with visual cues as a representative modality evaluated in real-world settings for robust deployment. Experiments in simulation and on real robots show that LG-Flow Policy achieves near single-step inference, improves trajectory smoothness by up to 93.7% and task success by up to 25% over raw action-space flow baselines, while remaining significantly faster than diffusion-based policies.
@@ I Introduction
Diffusion-based policies (DP) [6] have demonstrated strong performance in robotic visuomotor imitation learning, particularly for complex manipulation tasks in both simulation and real-world environments. By modeling multimodal action distributions conditioned on rich sensory observations such as images [6, 17] or point clouds [32, 25, 26], these methods enable expressive behavior modeling and generalization across diverse task variations.
@@
Despite these advances, deploying generative policies on physical robots remains challenging. Diffusion policies rely on iterative denoising, which introduces nontrivial inference latency and limits real-time applicability. Flow matching methods, while efficient, generate trajectories directly in the raw action space, where modeling noise and numerical errors tend to accumulate over long horizons, resulting in jittery trajectories and degraded execution stability. As a result, existing approaches struggle to simultaneously achieve expressive behavior modeling, fast inference, and stable execution.
Several recent works have explored different strategies to alleviate these issues. DP3 [32] and iDP3 [31] improve perception efficiency through lightweight point cloud encoders, RDP [28] introduces a hierarchical Fast–Slow control structure for long-horizon behaviors, and Flow Policy [34] incorporates consistency constraints to enhance execution performance. However, our experimental results indicate that the fundamental trade-off between inference efficiency and execution stability has not yet been fully resolved.
To address this challenge, we propose LG-Flow Policy, a trajectory-level action generation framework that performs flow matching in a continuous latent action space. By mapping action sequences into temporally regularized latent trajectory representations and learning an explicit latent-space flow, LG-Flow Policy preserves the inference efficiency of flow-based methods while substantially improving execution smoothness and stability. Operating on motion-level latent representations decouples global trajectory structure from low-level control noise, mitigating error amplification during trajectory generation and enabling more reliable long-horizon behaviors.
LG-Flow Policy further integrates geometry-aware point cloud conditioning and execution-time visual modulation from a wrist-mounted camera, providing rich environmental awareness without perturbing latent trajectory dynamics. Extensive simulation and real-world experiments demonstrate that the proposed approach consistently achieves smooth trajectories, high task success rates, and fast response times, outperforming both diffusion-based and raw action-space flow-based baselines.
@@
Figure 1: Overall architecture of the proposed LG-Flow Policy. The system first encodes point cloud observations into geometry-aware scene features, then performs latent-space flow matching to generate temporally coherent latent action trajectories, which are finally decoded into executable control commands under visual conditioning.
II Related Work
@@ II-A Diffusion and Flow-Based Policies for Robotics
Our work builds on consistency flow matching while addressing this limitation through trajectory-level latent modeling and geometry-aware conditioning.
@@ II-B Visual Imitation Learning with 3D Representations
Early visual imitation learning methods rely primarily on 2D image observations [6, 9], processed by convolutional networks [13] or vision transformers [8]. Recent 3D-based approaches [32, 31, 25, 12] incorporate depth or point clouds and improve geometric reasoning, but commonly face two challenges.
@@
II-C Trajectory-Level Modeling and Dexterous Manipulation
@@
Trajectory-level modeling improves long-horizon coherence and robustness to noisy or suboptimal demonstrations, with latent action representation playing a central role. Discrete latents via VQ-VAE [24] are widely adopted [1, 3, 4] to enforce information bottlenecks, but may limit expressiveness. Recent studies show that appropriately structured continuous latent representations can achieve strong performance without sacrificing stability [11, 15]. For example, CoMo [29] mitigates shortcut learning through temporal difference modeling, and LAOM [18] demonstrates the benefits of higher-capacity continuous bottlenecks.
These results motivate trajectory-level continuous latent action modeling as an effective mechanism for generating smooth and temporally coherent dexterous behaviors, which we adopt in this work.
@@ III-A Coherent Latent Action Representation
Long-horizon robotic manipulation requires action trajectories that are both expressive and execution-stable. Applying flow-based generative models directly in the raw action space often amplifies high-frequency noise during trajectory integration, leading to jittery and unreliable behaviors. To address this issue at its source, we introduce a trajectory-level continuous latent action representation that explicitly regularizes the geometry of the generation space before flow matching is applied. The overall architecture is illustrated in Fig. 2.
Trajectory-Level Latent Modeling
@@
The recurrent structure plays a critical role in shaping the latent space. By conditioning each latent code on the entire history of preceding action segments, the GRU introduces a strong temporal inductive bias that suppresses high-frequency variations and enforces trajectory-level coherence. As a result, the latent space forms a smooth trajectory manifold rather than a collection of independent step-level representations, which is particularly beneficial for stabilizing downstream ODE-based flow matching. Empirically, we observe that replacing the recurrent encoder with non-recurrent, per-chunk independent encoders leads to noticeably increased trajectory jitter.
Figure 2: Conditional variational autoencoder for latent action modeling. A GRU encoder enforces temporal coherence in the latent space, while an MLP decoder with FiLM-conditioned wrist-camera features enables visually adaptive action execution.
@@
To further regularize the latent action space, we adopt a variational formulation. The KL divergence between the posterior and a standard normal prior encourages compactness and continuity of the latent distribution, improving robustness to noisy or suboptimal demonstrations. Importantly, this regularization operates at the trajectory level rather than the control-step level, making it well suited for long-horizon manipulation.
@@
The proposed decoding mechanism is designed to flexibly incorporate execution-time sensory information without interfering with latent trajectory generation. In principle, when available, diverse modalities (e.g., vision, touch, and language) can be incorporated into the decoder through FiLM, enabling flexible multimodal conditioning while preserving coherent latent trajectory dynamics. Since the decoder operates at low computational cost, incorporating additional conditioning signals at this stage allows the policy to leverage richer information without sacrificing real-time responsiveness.
In this work, we instantiate this design using wrist-mounted camera observations in real-world experiments. Specifically, images captured at execution time are processed by a lightweight visual encoder, and the resulting features are used to modulate the decoder through FiLM layers. Given a latent action code $\mathbf{z}k$, the decoder reconstructs the corresponding action chunk as
$\hat{\mathbf{A}}^{(k)}=f\theta(\mathbf{z}_{k}\mid\mathbf{v}_{t}),$ (2)
where $\mathbf{v}_{t}$ denotes the encoded visual feature. By confining perceptual conditioning to the decoding stage, the system cleanly separates latent trajectory planning from execution-time adaptation, allowing low-level action realization to react to scene variations while maintaining the smoothness and temporal coherence of the latent plan.
Training Objective
The latent action representation is trained using a standard variational objective with a reconstruction term and a KL regularizer, augmented by a lightweight smoothness constraint between consecutive latent codes. While the smoothness term serves as an auxiliary regularizer, the primary source of temporal coherence arises from the recurrent encoding structure itself.
Overall, the proposed representation defines a low-dimensional, continuous, and temporally regularized latent action space. By reshaping the generation manifold prior to flow matching, it lays the foundation for stable and efficient long-horizon trajectory generation, which we describe next.
@@ III-B Latent-Space Action Generation via Flow Matching
Given a sequence of latent action codes $\{\mathbf{z}_{1},\ldots,\mathbf{z}_{K}\}$ obtained from the trajectory-level encoder, we generate long-horizon manipulation behaviors by modeling their evolution in continuous time using flow matching. Unlike diffusion-based policies that rely on iterative denoising, flow matching enables efficient one-step generation. However, when applied directly in the raw action space, flow-based models are highly sensitive to local inconsistencies and noise. We therefore perform flow matching exclusively in the regularized latent action space, which is explicitly designed to be smooth and temporally coherent (Sec. III-A).
@@
To efficiently learn the latent vector field across noise levels, we adopt consistency flow matching (CFM) and define the flow function
@@
to balance latent state scales across the time horizon.
@@
In contrast to diffusion-based policies that require iterative denoising through multiple inference steps, our approach learns an explicit latent-space flow field during training and therefore performs single-step integration at test time. This enables one-shot generation of the entire latent action trajectory without iterative sampling or numerical integration loops, substantially reducing inference latency.
The resulting latent trajectory preserves temporal coherence by construction and is decoded into executable action segments using the visual-conditioned decoder described in Sec. III-A. By decoupling trajectory generation from execution-level adaptation, the policy achieves both fast inference and stable, smooth manipulation behavior.
@@ III-C Geometry-Aware 3D Scene Conditioning
Stable execution of latent action trajectories requires accurate awareness of scene geometry, as geometric errors can propagate across multiple decoded action segments over long horizons. We therefore condition the latent flow policy on a compact, geometry-aware 3D scene representation derived from point clouds.
@@
Scene point clouds are reconstructed from depth images captured by a fixed global camera, followed by workspace cropping to remove irrelevant background. Farthest Point Sampling (FPS) is applied to select a sparse set of center points that uniformly cover the workspace, serving as anchors for local neighborhood construction.
@@
The extracted geometric features are injected into the latent flow network through a two-stage feature-wise linear modulation (FiLM) mechanism. Given an intermediate activation $\mathbf{h}$, local geometric modulation is applied first, followed by center-level conditioning:
@@ IV-A Evaluation Metrics
Task success rate is defined as the percentage of trials that satisfy task-specific success criteria within the episode horizon. Success is determined by predefined geometric constraints, including object pose errors below threshold and stable grasps or placements maintained for a minimum duration.
@@
$\mathcal{S}_{\mathrm{jerk}}=\frac{1}{T-2}\sum_{t=3}^{T}\|\mathbf{j}_{t}\|_{2}^{2}.$ (11)
@@
TABLE I: Main results for simulation experiments. Success rates (%) are reported as mean $\pm$ standard deviation. Avg. Success rate averages results over Adroit and MetaWorld. $\Delta$ Success / $\Delta$ Time denote relative changes w.r.t. DP3 (Baseline) in average success rate and per-step inference latency.
@@
LG-Flow Policy (Ours) $\mathbf{76.0\pm 2.0}$ $\mathbf{80.7\pm 3.1}$ $\mathbf{78.3\pm 3.4}$ $\mathbf{+12.6}$ $7.5\text{ ms}$ 48.4-48.4
@@ IV-B1 Simulation Setup
We compare our method against representative diffusion- and flow-based visuomotor policies. DP3 [32] serves as the primary point-cloud-based diffusion baseline, with iDP3 [31] and Reactive Diffusion Policy (RDP [28]) included as advanced variants. Due to environment constraints, RDP is evaluated without force or pressure sensing, and employs a standard VAE as a latent-space encoder on top of DP. We further include a 2D RGB-based Diffusion Policy [6] and a standard Flow Policy [34] to contrast 3D conditioning and latent-space modeling.
@@ IV-B2 Simulation Results Analysis
Simulation results are evaluated from three perspectives: trajectory smoothness, inference latency and task success.
@@
Fig. 4 reports trajectory smoothness across all simulated tasks. LG-Flow Policy achieves the lowest average smoothness score of $0.052$, reducing the metric by $51.4\%$, $69.4\%$, and $77.2\%$ relative to RDP, DP3, and Flow Policy, respectively. RDP exhibits lower smoothness than other diffusion-based baselines, indicating that latent action representations partially improve trajectory stability. However, its trajectories remain less smooth than those of LG-Flow Policy, as RDP does not impose explicit temporal smoothness constraints in the latent space. By contrast, LG-Flow Policy performs flow matching over temporally coherent latent trajectories, consistently suppressing high-frequency oscillations.
@@
Table I summarizes task success rates and inference latency across simulated benchmarks. LG-Flow Policy achieves the highest average success rate of $78.3\%$, outperforming Flow Policy ($17.3\%$) and DP3 ($12.6\%$), and improving over RDP ($73.0\%$) by $5.3\%$. While Flow Policy attains the lowest latency (6.1 ms), LG-Flow Policy incurs only a marginal increase (7.5 ms) and remains in the same real-time regime. Notably, LG-Flow Policy is approximately $7.5\times$ faster than DP3 (55.9 ms), demonstrating a favorable success–latency trade-off with minimal loss in responsiveness.
Overall, simulation results show that LG-Flow Policy achieves a favorable trade-off among smooth execution, fast inference, and high task success, which is essential for real-time robotic manipulation.

### v2

# CoLA-Flow Policy: Temporally Coherent Imitation Learning via Continuous Latent Action Flow Matching for Robotic Manipulation

@@ Abstract
Learning long-horizon robotic manipulation requires jointly achieving expressive behavior modelling, real-time inference, and stable execution, which remains challenging for existing generative policies. Diffusion-based approaches offer strong modelling capacity but incur high inference latency, while flow matching enables fast one-step generation yet often suffers from unstable execution when operating directly in the raw action space. We propose CoLA-Flow Policy, a trajectory-level imitation learning framework that performs flow matching in a continuous latent action space. By encoding action sequences into temporally coherent latent trajectories and learning an explicit latent-space flow, CoLA-Flow Policy decouples global motion structure from low-level control noise, enabling smooth and reliable long-horizon execution. The framework further integrates geometry-aware point cloud conditioning and execution-time multimodal modulation, with visual cues as a representative modality evaluated in real-world settings for robust deployment. Experiments in simulation and on real robots show that CoLA-Flow Policy achieves near single-step inference, improves trajectory smoothness by up to 93.7% and task success by up to 25% over raw action-space flow baselines, while remaining significantly faster than diffusion-based policies.
@@ I INTRODUCTION
Diffusion-based policies (DP) [5] have demonstrated strong performance in robotic visuomotor imitation learning, particularly for complex manipulation tasks in both simulation and real-world environments. By modelling multimodal action distributions conditioned on rich sensory observations such as images [5, 16] or point clouds [33, 25, 26], these methods enable expressive behavior modelling and generalization across diverse task variations.
@@
Despite these advances, deploying generative policies on physical robots remains challenging. Diffusion policies incur high inference latency due to iterative denoising, while flow matching methods generate actions by integrating a predicted flow field, a process that amplifies modelling noise over long horizons and leads to jittery, unstable trajectories. As a result, existing approaches struggle to simultaneously achieve expressive behavior modelling, fast inference, and stable execution.
Several recent works have explored different strategies to alleviate these issues. DP3 [33] and iDP3 [32] improve perception efficiency through lightweight point cloud encoders, RDP [28] introduces a hierarchical Fast–Slow control structure for long-horizon behaviours, and Flow Policy [35] incorporates consistency constraints to enhance execution performance. However, our experimental results indicate that the fundamental trade-off between inference efficiency and execution stability has not yet been fully resolved.
To address this challenge, we propose CoLA-Flow Policy, a trajectory-level action generation framework that performs flow matching in a continuous latent action space under geometry-aware 3D scene conditioning. Instead of generating controls directly in the physical action space, our approach models and generates entire action trajectories in a structured latent space, enabling efficient one-step inference while maintaining temporal coherence for long-horizon manipulation.
The key advantage of CoLA-Flow Policy lies in decoupling global motion structure from low-level control noise. By operating on motion-level latent representations, the learned flow captures smooth and consistent trajectory evolution, reducing the amplification of modelling errors during generation and leading to more stable execution. In addition, the latent trajectories can be decoded under environment-dependent modulation, allowing the policy to incorporate multi-modal sensory cues—such as visual feedback from a wrist-mounted camera at execution time—without altering the underlying generative process. This design results in reliable long-horizon behaviors with improved smoothness and robustness.
@@
Figure 1: Overall architecture of the proposed CoLA-Flow Policy. The system first encodes point cloud observations into geometry-aware scene features, then performs latent-space flow matching to generate temporally coherent latent action trajectories, which are finally decoded into executable control commands under visual conditioning.
II RELATED WORKS
@@ II-A Diffusion and Flow-Based Policies for Robotics
Our work builds on consistency flow matching while addressing this limitation through trajectory-level latent modelling and geometry-aware conditioning.
@@ II-B Visual Imitation Learning with 3D Representations
Early visual imitation learning methods rely primarily on 2D image observations [5, 8], processed by convolutional networks [12] or vision transformers [7]. Recent 3D-based approaches [33, 32, 25] incorporate depth or point clouds and improve geometric reasoning, but commonly face two challenges.
@@
II-C Trajectory-Level Modelling and Dexterous Manipulation
@@
Trajectory-level modelling improves long-horizon coherence and robustness to noisy or suboptimal demonstrations, with latent action representation playing a central role. Discrete latents via VQ-VAE [24] are widely adopted [2, 3] to enforce information bottlenecks, but may limit expressiveness. Recent studies show that appropriately structured continuous latent representations can achieve strong performance without sacrificing stability [10]. For example, CoMo [29] mitigates shortcut learning through temporal difference modelling, and LAOM [17] demonstrates the benefits of higher-capacity continuous bottlenecks.
These results motivate trajectory-level continuous latent action modelling as an effective mechanism for generating smooth and temporally coherent dexterous behaviours, which we adopt in this work.
@@ III METHODS
We propose CoLA-Flow Policy, a modular imitation learning framework for long-horizon robotic manipulation. As illustrated in Fig. 1, the method first learns a temporally coherent latent action representation, which serves as a compact trajectory-level abstraction. A generative model then operates in this latent space to produce long-horizon latent trajectories under geometric scene conditioning, and the resulting latent actions are finally decoded into executable control commands with execution-time sensory modulation.
@@ III-A Coherent Latent Action Representation
Long-horizon robotic manipulation requires action trajectories that are both expressive and execution-stable. Applying flow-based generative models directly in the raw action space often leads to numerical instability and high-frequency noise amplification during trajectory integration. We therefore introduce a trajectory-level latent action modeling framework that encodes action sequences into a continuous and temporally coherent latent trajectory space and decodes generated latent trajectories into executable controls, thereby reshaping the generation space prior to flow-based generative modeling. The overall architecture is illustrated in Fig. 2.
Temporally Continuous Latent Action Modeling
@@
By conditioning each latent code on the entire history of preceding action segments, the recurrent encoder induces a continuous and history-dependent latent trajectory, introducing a strong temporal inductive bias that suppresses high-frequency variations. As a result, the latent action space is shaped into a smooth and temporally continuous trajectory manifold rather than a collection of independent per-chunk representations, which is critical for stable ODE-based flow matching. In contrast, non-recurrent per-chunk encoders break this continuity and empirically lead to increased trajectory jitter.
Figure 2: Trajectory-level latent action representation with recurrent encoding and conditional decoding. A GRU-based encoder introduces a temporal inductive bias that promotes coherent latent trajectories, while an MLP decoder modulated by wrist-camera features via FiLM enables visually adaptive action execution.
@@
To further regularize the latent space, we adopt a variational formulation [14]. A KL divergence term between the posterior and a standard normal prior encourages compactness and continuity of the latent distribution, improving robustness to noisy demonstrations. This regularization operates at the trajectory level, making it well suited for long-horizon manipulation.
@@
The decoder is designed to incorporate execution-time sensory information without interfering with latent trajectory generation. In principle, diverse modalities(e.g., vision, tactile, proprioception) can be integrated through feature-wise linear modulation (FiLM) [18], enabling flexible multimodal conditioning at low computational cost.
In this work, we instantiate this design using wrist-mounted camera observations, which are encoded by a lightweight visual encoder based on a pretrained ResNet-18. [11]. Execution-time images are processed by a lightweight visual encoder and injected into the decoder via FiLM layers. Given a latent action code $\mathbf{z}_{k}$, the decoder reconstructs the corresponding action chunk as
$\hat{\mathbf{A}}^{(k)}=f_{\theta}(\mathbf{z}_{k}\mid\mathbf{v}_{t}),$ (2)
where $\mathbf{v}_{t}$ denotes the encoded visual feature. By confining perceptual conditioning to the decoding stage, the framework cleanly separates latent trajectory planning from execution-time adaptation while preserving temporal coherence.
Training Objective and Protocol
The latent action encoder–decoder is first trained using a variational objective with reconstruction, KL regularization, and a lightweight smoothness constraint to learn a temporally coherent latent space. The learned representation is then fixed and used as a plug-in interface, upon which a downstream generative model is trained to generate long-horizon latent trajectories.
@@ III-B Latent-Space Action Generation via Flow Matching
Given a sequence of latent action codes $\{\mathbf{z}_{1},\ldots,\mathbf{z}_{K}\}$ obtained from the trajectory-level encoder, we generate long-horizon manipulation behaviors by modelling their evolution in continuous time using flow matching [15]. Unlike diffusion-based policies [5] that rely on iterative denoising, flow matching enables efficient one-step generation. However, when applied directly in the raw action space, flow-based models are highly sensitive to local inconsistencies and noise. We therefore perform flow matching exclusively in the temporally coherent latent action space, which is explicitly designed to be smooth and temporally coherent (Sec. III-A).
@@
To efficiently learn the latent vector field across noise levels, we adopt consistency flow matching (CFM) [30] and define the flow function
@@
scaling the latent state before feeding it into the velocity field network: $\boldsymbol{\nu}_{\theta}\!\left(t,\,c_{\mathrm{in}}(t)\cdot\mathbf{z}\right)$. This normalization balances latent state scales across noise levels and prevents overfitting to specific time regions.
@@
As with flow-based generative models, latent trajectories are obtained through a single-step transformation defined by the learned latent-space flow field, without requiring iterative denoising as in diffusion-based policies [5]. This enables one-shot generation of the entire latent action trajectory, avoiding iterative sampling and substantially reducing inference latency.
The resulting latent trajectory preserves temporal coherence by construction and is decoded into executable action segments using the multimodal-conditioned decoder described in Sec. III-A. By decoupling trajectory generation from execution-level adaptation, the policy achieves both fast inference and stable, smooth manipulation behavior.
@@ III-C Geometry-Aware 3D Scene Conditioning
Stable execution of latent action trajectories requires accurate awareness of scene geometry, as geometric errors can propagate across multiple decoded action segments over long horizons. We therefore condition the latent flow policy on a compact, geometry-aware 3D scene representation derived from point clouds, following prior point-cloud conditioned manipulation policies [33].
@@
Scene point clouds are reconstructed from depth images captured by a fixed global camera, followed by workspace cropping to remove irrelevant background. Farthest Point Sampling (FPS) [19] is applied to select a sparse set of center points that uniformly cover the workspace, serving as anchors for local neighborhood construction.
@@
The extracted geometric features are injected into the latent flow network through a two-stage FiLM mechanism. Given an intermediate activation $\mathbf{h}$, local geometric modulation is applied first, followed by center-level conditioning:
@@ IV-A Evaluation Metrics
Task success rate is defined as the percentage of trials that satisfy task-specific success criteria within the episode horizon. Success is determined by predefined geometric constraints, including object pose errors below a threshold and stable grasps or placements maintained for a minimum duration.
@@
$\mathcal{S}_{\mathrm{jerk}}=\frac{1}{T-2}\sum\nolimits_{t=3}^{T}\|\mathbf{j}_{t}\|_{2}^{2}.$ (11)
@@
TABLE I: Main results for simulation experiments. Success rates (%) are reported as mean $\pm$ standard deviation. Avg. Success rate averages results over Adroit and MetaWorld. $\Delta$ Success / $\Delta$ Time denotes relative changes w.r.t. DP3 in average success rate and per-step inference latency.
@@
CoLA-Flow Policy (Ours) $\mathbf{76.0\pm 2.0}$ $\mathbf{80.7\pm 3.1}$ $\mathbf{78.3\pm 3.4}$ $\mathbf{+12.6}$ $7.5\text{ ms}$ 48.4-48.4
@@ IV-B1 Simulation Setup
We compare our method against representative diffusion- and flow-based visuomotor policies. DP3 [33] serves as the primary point-cloud-based diffusion baseline, with iDP3 [32] and Reactive Diffusion Policy (RDP [28]) included as advanced variants, where the former employs a more expressive scene encoder, and the latter augments diffusion-based policies with an additional non-recurrent encoder. Due to environment constraints, RDP is evaluated without force or pressure sensing. We further include the original 2D RGB-based Diffusion Policy [5] as a canonical reference in visuomotor imitation learning, and a standard Flow Policy [35] that applies continuous flow matching directly to robotic action generation, serving as a representative flow-based control baseline.
@@ IV-B2 Simulation Results Analysis
Quantitative real-world results are summarized in Table I and Fig. 4, evaluating trajectory smoothness, task success, and response time.
@@
Fig. 4 reports trajectory smoothness across all simulated tasks. CoLA-Flow Policy achieves the lowest average smoothness score of $0.052$, reducing the metric by $51.4\%$, $69.4\%$, and $77.2\%$ relative to RDP, DP3, and Flow Policy, respectively. RDP exhibits lower smoothness than other diffusion-based baselines, indicating that latent action representations partially improve trajectory stability. However, its trajectories remain less smooth than those of CoLA-Flow Policy, as RDP does not adopt temporally continuous latent action modeling and does not explicitly structure its latent space to be continuous over time. By contrast, CoLA-Flow Policy performs flow matching over temporally coherent latent trajectories, consistently suppressing high-frequency oscillations.
@@
Table I summarizes task success rates and inference latency across simulated benchmarks. CoLA-Flow Policy achieves the highest average success rate of $78.3\%$, outperforming Flow Policy ($17.3\%$) and DP3 ($12.6\%$), and improving over RDP ($73.0\%$) by $5.3\%$. While Flow Policy attains the lowest latency (6.1 ms), CoLA-Flow Policy incurs only a marginal increase (7.5 ms) and remains in the same real-time regime. Notably, CoLA-Flow Policy is approximately $7.5\times$ faster than DP3 (55.9 ms), demonstrating a favorable success–latency trade-off with minimal loss in responsiveness.
Overall, simulation results show that CoLA-Flow Policy achieves a favorable trade-off among smooth execution, fast inference, and high task success, which is essential for real-time robotic manipulation.